OpenStudy (anonymous):

What is the equation of a circle with a center at (4, 0) and a radius of 1 A (x – 4)2 + y2 = 1 B (x + 4)2 + y2 = 1 C x2 + (y – 4)2 = 1 D (x – 4)2 + y2 = 2

OpenStudy (anonymous):

@julie001 can you do a step by step. please

OpenStudy (anonymous):

*(x-h)^2 + (y-k)^2 = r^2 the center is already given as well as the radius. I wont do it for u but i'll help u each step.

OpenStudy (anonymous):

Where the h & the k is u substitute that for the center (4,0) and radius being 1 which is where the r is

OpenStudy (anonymous):

@julie001 (x + 4)2 + y2 = 1

OpenStudy (anonymous):

no u have to follow the symbols in the equation

OpenStudy (anonymous):

@julie so it would be (-)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

basically @julie001 (x – 4)2 + y2 = 1

OpenStudy (anonymous):

u got it !

OpenStudy (anonymous):

@julie001 thank you!!! can you double check my answers for the next 3 real quicK?

OpenStudy (anonymous):

lets see them

OpenStudy (anonymous):

What is the equation of a circle with a center at (0, 9) and a radius of 9? x2 + (y – 9)2 = 81

OpenStudy (anonymous):

@julie001 thats what i got

OpenStudy (anonymous):

u got it!

OpenStudy (anonymous):

@julie001 cool. What is the equation of a circle with a center at (0, 0) and a radius of 4.5? i got (x2 + y2 = 20.25)

OpenStudy (anonymous):

yep, and i didnt have to work it out

OpenStudy (anonymous):

@julie001 yes! What is the equation of a circle with a center at (-4, -8) and a radius of 1.9? i got (x + 4)2 + (y + 8)2 = 3.61

OpenStudy (anonymous):

ur absolutely right! u didnt even need my help u got this :)

OpenStudy (anonymous):

@julie001 i did need you!!! thank you so much.. i have to move on. thanks again.

OpenStudy (anonymous):

yw